# Multi-label Guided Soft Contrastive Learning for Efficient Earth Observation Pretraining

### v1

@@ Abstract
Self-supervised pretraining on large-scale satellite data has raised great interest in building Earth observation (EO) foundation models. However, many important resources beyond pure satellite imagery, such as land-cover-land-use products that provide free global semantic information, as well as vision foundation models that hold strong knowledge of the natural world, tend to be overlooked. In this work, we show these free additional resources not only help resolve common contrastive learning bottlenecks, but also significantly boost the efficiency and effectiveness of EO pretraining.
Specifically, we first propose soft contrastive learning that optimizes cross-scene soft similarity based on land-cover-generated multi-label supervision, naturally solving the issue of multiple positive samples and too strict positive matching in complex scenes. Second, we explore cross-domain continual pretraining for both multispectral and SAR imagery, building efficient EO foundation models from strongest vision models such as DINOv2. Integrating simple weight-initialization and Siamese masking strategies into our soft contrastive learning framework, we demonstrate impressive continual pretraining performance even when the input channels and modalities are not aligned.
Without prohibitive training, we produce multispectral and SAR foundation models that achieve significantly better results in 9 out of 10 downstream tasks than most existing SOTA models. For example, our ResNet50/ViT-S achieve 84.8/85.0 linear probing mAP scores on BigEarthNet-10% which are better than most existing ViT-L models; under the same setting, our ViT-B sets a new record of 86.8 in multispectral, and 82.5 in SAR, the latter even better than many multispectral models. Dataset and models are available at https://github.com/zhu-xlab/softcon.
@@ I Introduction
Self-supervised learning has driven wide attention in pretraining Earth observation (EO) foundation models on large-scale satellite data [1, 2, 3]. While more and more efforts are spent on scaling up the data and model size with purely unsupervised pretraining, many other resources such as various land cover land use products tend to be overlooked. For example, ESA WorldCover [4, 5] provides the first global land cover maps for 2020 and 2021 at 10 m resolution, and Google Dynamic World [6] provides a continuous dataset of near-real-time land use land cover mapping. These dense products are highly correlated with commonly studied medium-resolution satellite imagery, and offer free semantic annotations with real-global coverage. Even though they are noisy at pixel-level due to semi-automatic creation process, they can be easily integrated into scene-level annotations with rather good quality. In this work, we will show the benefits of such auxiliary information in extending the popular contrastive self-supervised learning framework to build EO foundation models.
@@
Meanwhile, another important resource that has huge potential in helping build EO foundation models lies in the general computer vision community: the vision foundation models. These models are exhaustively trained on huge amount of natural images, and have already gained strong knowledge of the visual world. Dating back to before the era of foundation models, ImageNet pretrained weights had been widely used and proved beneficial in many supervised Earth observation tasks. Similarly, they can also be used in unsupervised learning, leading to cross-domain continual pretraining. In this regard, recent works such as GFM [11] propose to build EO foundation models by distilling frozen ImageNet models. However, GFM-style training is limited to RGB images, restricting the flexibility to adapt to various EO sensors. Furthermore, the fast advances in computer vision have made available much stronger vision models than ImageNet supervised weights. To bridge this gap, we revisit and verify the advantage of cross-domain continual pretraining with simple weight initialization. For unaligned input modalities, we simply leave the first layer’s weights randomly initialized. In addition, to save hardware memory when continually training large Vision Transformers, we adopt Siamese masking inspired by masked Autoencoder [12]. Specifically, we randomly mask out a certain percentage of input patches on the trainable branch of the Siamese contrastive learning framework, and only send the remaining visible patches to the encoder. We show such simple but flexible continual pretraining strategies, when applied with strong vision foundation models such as DINOv2 [13], exhibit impressive effectiveness and efficiency even when the imaging sensors are completely different from the source domain.
Integrating the continual pretraining strategies into the soft contrastive learning framework, we efficiently train CNN and ViT foundation models that reach SOTA performances in 9 out of 10 downstream tasks. For example, our ResNet50/ViT-S (23M parameters, 100 epochs) achieve 84.8/85.0 linear probing mAP scores on BigEarthNet-10% which are better than most existing ViT-L models (300M parameters, 100-300 epochs) and comparable to CROMA and SkySense ($\geq$600 epochs); under the same setting, our ViT-B (86M parameters) sets a new SOTA of 86.8 in multispectral, and 82.5 in SAR, the latter even better than most existing multispectral models.
@@
  • We explore the benefits of open resources beyond pure satellite imagery for efficient EO pretraining, producing multispectral and SAR foundation models that reach SOTA performances in 9 out of 10 downstream tasks.
@@
  • We propose cross-domain continual pretraining with simple weight initialization from strong vision foundation models and Siamese masking. We verify that even when the input modalities are not aligned, the strong knowledge can still be efficiently transferred to the target EO domain.
@@ II-A Earth observation foundation models
Massive research has been conducted on the development of Earth observation (EO) foundation models, a majority of which tackle the technical adaptation of self-supervised pretraining techniques into EO domain. Early works focus on EO-specific data characteristics for contrastive view generation. For example, Tile2vec [14] proposed to pull together geospatially close tiles while pushing away far tiles. SeCo [15] proposed to use different seasons as augmented views for contrastive learning. CACo [16] proposed to perceive temporal changes with the spatiotemporal structure of remote sensing time series. Further, another group of works explores masked-image-modeling-based pretraining [17]. SatMAE [18] proposed temporal and spectral positional encoding for multispectral imagery and time series. RingMo [19] proposed less aggressive masking. Scale-MAE [20] proposed GSD-based positional encoding and multi-scale reconstruction. SpectralGPT [21] proposed 3D masking to encode and reconstruct spectral data. FG-MAE [22] proposed to reconstruct remote sensing image features such as normalized difference indices. Going beyond a single modality and time stamp, CROMA [23], DeCUR [24], OFA-Net [25] and DOFA [26] investigated multi-modal pre-training, RingMo-sense [27], Presto [28] and Prithvi [29] studied EO time series, and SkySense [30] combined both modality and time sequence in a unified architecture, reaching SOTA performance in many downstream tasks. Meanwhile, another line of research targets the curation of EO pretraining datasets, such as SSL4EO-S12 [9] for Sentinel-1 and 2, SSL4EO-L [31] for Landsat series, and SatlasPretrain [32] for medium- and high-resolution satellite and aerial imagery with extensive annotations. While a general trend is to scale the data and model sizes with exhaustive training cost, important existing resources such as open annotations and vision foundation models tend to be overlooked. In this study, we propose to use land-cover-generated multilabel annotations to guide EO pretraining, and share the insight with GFM [11] and DOFA to utilize vision models for continual pretraining.
@@ II-B Multi-positive contrastive learning
Contrastive learning beyond simple instance discrimination has been widely explored in computer vision. SupCon [10] proposed to use image labels for positive matching, extending contrastive learning to the fully-supervised setting. Almost at the same time, ML-CPC [33] proposed multi-label contrastive predictive coding, identifying multiple positive samples as a multi-label classification problem [^1] We note that depending on the context, ”multi-label” in this paper may refer to two things: 1) the contrastive learning objective with multi-positive matching; 2) the input images are multi-label annotated.. Further on, WCL [34] combined contrastive instance discrimination with SupCon by predicting pseudo weak labels. Sel-CL [35] extended SupCon to deal with noisy labels. All of these methods still deal with single-label datasets. To deal with more complex scenes, HiMulCon [36] presented a hierarchical representation learning framework that can leverage all available labels and preserve the hierarchical relationship between classes. MLS [37] proposed to assign multiple binary pseudo-labels for each input image by comparing its embeddings with those in two dictionaries, and training the model with binary cross-entropy loss. However, these two methods are still based on hard multi-positive matching, restricting the soft relationship between images with overlapped multi-label distributions. To tackle this issue, and to make the best use of real-world multi-label annotations, we propose soft contrastive learning (SoftCon) that allows soft positive matching between different multi-label scenes.
@@ II-C Continual pretraining
Early works of continual pretraining were mainly developed in natural language processing to improve domain-specific large language models [38, 39]. In vision, existing works like [40] and [41] proposed hierarchical pretraining approaches for task adaptation, while not targeting task-agnostic representations. Recently, SpectralGPT [21] used a progressive pretraining pipeline between EO datasets to benefit from their unique advantages, yet it still requires first-stage pretraining from scratch. GFM [11] and DOFA [26] explored continual pretraining for generic EO foundation models by distilling knowledge from frozen vision backbones. However, it heavily relies on ImageNet weights whose prior knowledge is restricted compared to stronger vision foundation models. Meanwhile, GFM-style continual pretraining can only process RGB data, limiting the flexibility to adapt to various EO modalities such as multispectral imagery. To bridge this gap, we propose to revisit the simple cross-domain continual pretraining with weight initialization from strong vision foundation models such as DINOv2 [13]. For non-RGB imagery, we simply leave the input embedding layer randomly initialized. We show such a strategy, though naive, is both flexible and impressively effective in EO pretraining regardless of target sensors.
@@ III-A Building a large-scale multi-label dataset
Figure 3: Different contrastive learning designs. (a) The original contrastive learning performs strict instance discrimination where one anchor image has only one positive pair; (b) supervised contrastive learning allows for multiple positive responses when images belong to the same class; (c) our proposed soft contrastive learning can effectively exploit multi-label annotations by assigning soft similarity scores to different pairs. SofCon is a more generic design that covers SupCon: when multi-label degrades to single-label, the soft similarities turn into binary scores, thus becoming multi-positive supervised contrastive learning.
@@ III-B Multi-label guided soft contrastive learning
Figure 4: The general framework of SoftCon. We load vision foundation models for both the trainable and the momentum encoder when initializing the model. EMA represents exponential moving average. Siamese masking is only used with ViT backbones.
@@
where $\tau$ is the softmax temperature. In practice, contrastive learning typically needs a large number of negative samples. Therefore, it inevitably faces the conflict that multiple images besides only the augmented view can belong to the same class as the anchor image, but the InfoNCE loss strictly pushes them apart.
@@
To effectively exploit multi-label annotations, we propose soft contrastive learning as is shown in Figure 3 (c). Samples with more similar label distributions are pulled closer than samples with more dissimilar labels, resulting in a soft positive matching. Specifically, we calculate the similarity matrix $Y\in\mathbb{R}^{N\times N}$ of the normalized multi-hot label vectors $\{y_{1},y_{2},...,y_{N}\}$ and $\{y^{\prime}_{1},y^{\prime}_{2},...,y^{\prime}_{N}\}$. Similarly, we get the similarity matrix $X\in\mathbb{R}^{N\times N}$ of the feature vectors $\{z_{1},z_{2},...,z_{N}\}$ and $\{z^{\prime}_{1},z^{\prime}_{2},...,z^{\prime}_{N}\}$. Then, we optimize per-element binary cross entropy loss:
@@ III-C Continual pretraining with Siamese masking
Finally, we introduce cross-domain continual pretraining into our framework to boost efficient Earth observation (EO) pretraining. Instead of sophisticated strategies like GFM [11] that are restricted by RGB input, we revisit the simple weight loading, initializing the backbone with strong vision foundation models. We use DINO [49] weights for ResNet backbones, and DINOv2 [13] weights for ViTs. As for the channel difference between natural RGB images and EO multispectral and SAR imagery, we simply let the input embedding layer remain randomly initialized. Our experiments empirically suggest that though naive, this strategy is both flexible and effective. In addition, to save hardware memory when continually training large Vision Transformers, we adopt Siamese masking inspired by iBOT [50], MAE [12] and MSN [51]. Specifically, we randomly mask out a certain percentage of input patches on the trainable branch of the Siamese contrastive learning framework, and only send the remaining visible patches to the encoder. Note that we do not conduct any reconstruction, but rather view such masking as additional data augmentation. Different from MSN’s non-negative online clustering, we verify it also works well in contrastive learning with negative sampling: with a reasonable masking ratio of about 20%, not only the efficiency, but also the performance can be improved. The full pretraining framework of SoftCon is illustrated in Figure 4.
@@ IV-A Pretraining
We load ResNet50 weights from DINO [^2] https://github.com/facebookresearch/dino [49] and ViT-S/14 and ViT-B/14 weights (without register) from DINOv2333https://github.com/facebookresearch/dinov2 [13], and conduct continual pretraining for 100 epochs. For ViTs, we randomly mask out 20% patches and send the remaining ones to the trainable encoder. More details can be seen in the appendix.

### v2

@@ Abstract
Self-supervised pretraining on large-scale satellite data has raised great interest in building Earth observation (EO) foundation models. However, many important resources beyond pure satellite imagery, such as land-cover-land-use products that provide free global semantic information, as well as vision foundation models that hold strong knowledge of the natural world, are not widely studied. In this work, we show these free additional resources not only help resolve common contrastive learning bottlenecks, but also significantly boost the efficiency and effectiveness of EO pretraining.
Specifically, we first propose soft contrastive learning that optimizes cross-scene soft similarity based on land-cover-generated multi-label supervision, naturally solving the issue of multiple positive samples and too strict positive matching in complex scenes. Second, we revisit and explore cross-domain continual pretraining for both multispectral and SAR imagery, building efficient EO foundation models from strongest vision models such as DINOv2. Adapting simple weight-initialization and Siamese masking strategies into our soft contrastive learning framework, we demonstrate impressive continual pretraining performance even when the input modalities are not aligned.
Without prohibitive training, we produce multispectral and SAR foundation models that achieve significantly better results in 10 out of 11 downstream tasks than most existing SOTA models. For example, our ResNet50/ViT-S achieve 84.8/85.0 linear probing mAP scores on BigEarthNet-10% which are better than most existing ViT-L models; under the same setting, our ViT-B sets a new record of 86.8 in multispectral, and 82.5 in SAR, the latter even better than many multispectral models. Dataset and models are available at https://github.com/zhu-xlab/softcon.
@@ I Introduction
Self-supervised learning has driven wide attention in pretraining Earth observation (EO) foundation models on large-scale satellite data [1, 2, 3]. While more and more efforts are spent on scaling up the data and model size with purely unsupervised pretraining, many other resources such as various land cover land use products tend to be overlooked. For example, ESA WorldCover [4, 5] provides the first global land cover maps for 2020 and 2021 at 10 m resolution, and Google Dynamic World [6] provides a continuous dataset of near-real-time land use land cover mapping. These dense products are highly correlated with commonly studied medium-resolution satellite imagery, and offer free semantic annotations with real-global coverage. Even though they are noisy at pixel-level due to semi-automatic creation process, they can be easily integrated into scene-level annotations with rather good quality. A similar concept has been demonstrated valid in supervised pretraining in GeoKR [7], where land cover products and geographical location are regarded as geographical knowledge to provide supervision. In this work, we will show the benefits of such auxiliary information in extending the popular contrastive self-supervised learning framework to build EO foundation models.
@@
Meanwhile, another important resource that has huge potential in helping build EO foundation models lies in the general computer vision community: the vision foundation models. These models are exhaustively trained on huge amount of natural images, and have already gained strong knowledge of the visual world. Dating back to before the era of foundation models, ImageNet pretrained weights had been widely used and proved beneficial in many supervised Earth observation tasks. Similarly, they can also be used in unsupervised learning, leading to cross-domain continual pretraining. In this regard, recent works such as GFM [12] propose to build EO foundation models by distilling frozen ImageNet models. However, GFM-style training is limited to RGB images, restricting the flexibility to adapt to various EO sensors. Furthermore, the fast advances in computer vision have made available much stronger vision models than ImageNet supervised weights. To bridge this gap, we revisit the natural idea of simple weight initialization which has been preliminarily explored and verified effective in recent works [13, 14]. For unaligned input modalities, we simply leave the first layer’s weights randomly initialized. In addition, to save hardware memory when continually training large Vision Transformers, we adopt Siamese masking inspired by masked Autoencoder [15]. Specifically, we randomly mask out a certain percentage of input patches on the trainable branch of the Siamese contrastive learning framework, and only send the remaining visible patches to the encoder. We show such simple but flexible continual pretraining strategies, when applied with strong vision foundation models such as DINOv2 [16], exhibit impressive effectiveness and efficiency even when the imaging sensors are completely different from the source domain.
Integrating the continual pretraining strategies into the soft contrastive learning framework, we efficiently train CNN and ViT foundation models that reach SOTA performances in 10 out of 11 downstream tasks. For example, our ResNet50/ViT-S (23M parameters, 100 epochs) achieve 84.8/85.0 linear probing mAP scores on BigEarthNet-10% which are better than most existing ViT-L models (300M parameters, 100-300 epochs) and comparable to CROMA and SkySense ($\geq$600 epochs); under the same setting, our ViT-B (86M parameters) sets a new SOTA of 86.8 in multispectral, and 82.5 in SAR, the latter even better than most existing multispectral models.
@@
  • We explore the benefits of open resources beyond pure satellite imagery for efficient EO pretraining, producing multispectral and SAR foundation models that reach SOTA performances in 10 out of 11 downstream tasks.
@@
  • We revisit cross-domain continual pretraining with simple weight initialization from strong vision foundation models and Siamese masking. We verify that even when the input modalities are not aligned, the strong knowledge can still be efficiently transferred to the target EO domain.
@@ II-A Earth observation foundation models
Massive research has been conducted on the development of Earth observation (EO) foundation models. While there is also a line of supervised pretraining [17] on large-scale labeled datasets [18, 19], a majority of works tackle the technical adaptation of self-supervised pretraining techniques into EO domain. Early works focus on EO-specific data characteristics for contrastive view generation. For example, Tile2vec [20] proposed to pull together geospatially close tiles while pushing away far tiles. SeCo [21] proposed to use different seasons as augmented views for contrastive learning. CACo [22] proposed to perceive temporal changes with the spatiotemporal structure of remote sensing time series. Further, another group of works explores masked-image-modeling-based pretraining [23]. SatMAE [24] proposed temporal and spectral positional encoding for multispectral imagery and time series. RingMo [25] proposed less aggressive masking. Scale-MAE [26] proposed GSD-based positional encoding and multi-scale reconstruction. SpectralGPT [27] proposed 3D masking to encode and reconstruct spectral data. FG-MAE [28] proposed to reconstruct remote sensing image features such as normalized difference indices. Going beyond a single modality and time stamp, CROMA [29], DeCUR [30], OFA-Net [31] and DOFA [32] investigated multi-modal multi-sensor pre-training, RemoteCLIP [33], SkyScript [34] and BITA [35] explored EO vision language pretraining, RingMo-sense [36], Presto [37] and Prithvi [38] studied EO time series, and SkySense [39] combined both modality and time sequence in a unified architecture, reaching SOTA performance in many downstream tasks. Meanwhile, another line of research targets the curation of EO pretraining datasets, such as SSL4EO-S12 [10] for Sentinel-1 and 2, SSL4EO-L [40] for Landsat series, and SatlasPretrain [41] for medium- and high-resolution satellite and aerial imagery with extensive annotations. While a general trend is to scale the data and model sizes with exhaustive training cost, important existing resources such as open annotations and vision foundation models tend to be overlooked. In this study, we share the conceptual insight with GeoKR [7] to use land-cover-generated multilabel annotations to guide EO pretraining, and with GFM [12] and DOFA to utilize vision models for continual pretraining.
@@ II-B Multi-positive contrastive learning
Contrastive learning beyond simple instance discrimination has been widely explored in computer vision. SupCon [11] proposed to use image labels for positive matching, extending contrastive learning to the fully-supervised setting. Almost at the same time, ML-CPC [42] proposed multi-label contrastive predictive coding, identifying multiple positive samples as a multi-label classification problem [^1] We note that depending on the context, ”multi-label” in this paper may refer to two things: 1) the contrastive learning objective with multi-positive matching; 2) the input images are multi-label annotated.. Further on, WCL [43] combined contrastive instance discrimination with SupCon by predicting pseudo weak labels. Sel-CL [44] extended SupCon to deal with noisy labels. All of these methods still deal with single-label datasets. To deal with more complex scenes, HiMulCon [45] presented a hierarchical representation learning framework that can leverage all available labels and preserve the hierarchical relationship between classes. MLS [46] proposed to assign multiple binary pseudo-labels for each input image by comparing its embeddings with those in two dictionaries, and training the model with binary cross-entropy loss. However, these two methods are still based on hard multi-positive matching, restricting the soft relationship between images with overlapped multi-label distributions. To tackle this issue, and to make the best use of real-world multi-label annotations, we propose soft contrastive learning that allows soft positive matching between different multi-label scenes.
@@ II-C Continual pretraining
Early works of continual pretraining were mainly developed in natural language processing to improve domain-specific large language models [47, 48]. In vision, existing works like [49] and [50] proposed hierarchical pretraining approaches for task adaptation, while not targeting task-agnostic representations. In remote sensing, CSPT [51] and TOV [52] proposed consecutive pretraining from natural images to remote sensing images, yet they are both limited to retraining on natural images for the source model. SpectralGPT [27] used a similar progressive pretraining pipeline between EO datasets to benefit from their unique advantages, yet it still requires first-stage pretraining from scratch. GFM [12] and DOFA [32] explored continual pretraining for generic EO foundation models by distilling knowledge from frozen vision backbones. However, it heavily relies on ImageNet weights whose prior knowledge is restricted compared to stronger vision foundation models. Meanwhile, GFM-style continual pretraining can only process RGB data, limiting the flexibility to adapt to various EO modalities such as multispectral imagery. To bridge this gap, we revisit and explore the simple cross-domain continual pretraining with weight initialization [13, 14] from strong vision foundation models such as DINOv2 [16]. For non-RGB imagery, we simply leave the input embedding layer randomly initialized. We show such a strategy, though naive, is both flexible and impressively effective in EO pretraining regardless of target sensors.
@@ III-A Building a large-scale multi-label dataset
Figure 3: Different contrastive learning designs. (a) The original contrastive learning performs strict instance discrimination where one anchor image has only one positive pair; (b) supervised contrastive learning allows multiple positive responses when images belong to the same class; (c) our proposed soft contrastive learning can effectively exploit multi-label annotations by assigning soft similarity scores to different pairs. SofCon is a more generic design that covers SupCon: when multi-label degrades to single-label, the soft similarities turn into binary scores, thus becoming multi-positive supervised contrastive learning.
@@ III-B Multi-label guided soft contrastive learning
Figure 4: The general framework of SoftCon. Given a batch of input images, two batches of augmented views are parallelly sent through the two branches of a Siamese network. A similarity matrix is calculated based on the resulting two batches of feature vectors. A weighted sum of the contrastive and soft contrastive loss is optimized. For contrastive loss, this matrix should be close to Identity; for soft contrastive loss, this matrix should be close to the similarity matrix of the label vectors. We load vision foundation models for both the trainable (base encoder) and the frozen (momentum encoder) branches when initializing the model. During training, the weights of the momentum encoder are updated by exponential moving average (EMA) of the base encoder: $\theta_{\mathrm{momentum}}\leftarrow m\cdot\theta_{\mathrm{momentum}}+(1-m)\cdot\theta_{\mathrm{base}}$, where $m\in[0,1)$ is a momentum coefficient. Siamese masking is only used with ViT backbones, where random patches of an image are masked out and only the visible patches are sent through the trainable branch.
@@
where $\tau$ is softmax temperature. In practice, contrastive learning typically needs a large number of negative samples. Therefore, it inevitably faces the conflict that multiple images besides the augmented view can belong to the same class as the anchor image while the loss strictly pushes them apart.
@@
To effectively exploit multi-label annotations, we propose soft contrastive learning as is shown in Figure 3 (c). Samples with more similar label distributions are pulled closer than samples with more dissimilar labels, resulting in a soft positive matching. Specifically, we calculate the pair-wise cosine similarity matrix $Y\in\mathbb{R}^{N\times N}$ of the normalized multi-hot label vectors $\{y_{1},y_{2},...,y_{N}\}$ and $\{y^{\prime}_{1},y^{\prime}_{2},...,y^{\prime}_{N}\}$. This is done through the dot product of the label vectors: $\mathbf{y}\mathbf{{y^{\prime}}^{T}}$. Similarly, we get the similarity matrix $X=\mathbf{z}\mathbf{{z^{\prime}}^{T}}\in\mathbb{R}^{N\times N}$ of the feature vectors $\{z_{1},z_{2},...,z_{N}\}$ and $\{z^{\prime}_{1},z^{\prime}_{2},...,z^{\prime}_{N}\}$. Then, we optimize per-element binary cross entropy loss:
@@ III-C Continual pretraining with Siamese masking
Finally, we introduce cross-domain continual pretraining into our framework to boost efficient Earth observation (EO) pretraining. Instead of sophisticated strategies like GFM [12] that are restricted by RGB input, we revisit the simple weight loading, initializing the backbone with strong vision foundation models. We use DINO [60] weights for ResNet backbones, and DINOv2 [16] weights for ViTs. As for the channel difference between natural RGB images and EO multispectral and SAR imagery, we simply let the input embedding layer remain randomly initialized. Our experiments empirically suggest that though naive, this strategy is both flexible and effective. In addition, to save hardware memory when continually training large vision Transformers, we adopt Siamese masking inspired by iBOT [61], MAE [15] and MSN [62]. Specifically, we randomly mask out a certain percentage of input patches on the trainable branch of the Siamese contrastive learning framework, and only send the remaining visible patches to the encoder. As the encoder only needs to process a portion of the full patches, both the memory and the training time can be reduced. Note that we do not conduct any reconstruction like MAE or iBOT, but rather view such masking as additional data augmentation. Different from MSN’s non-negative online clustering, we verify it also works well in contrastive learning with negative sampling: with a reasonable masking ratio of about 20%, not only the efficiency, but also the performance can be improved. The full pretraining framework of SoftCon is illustrated in Figure 4.
@@ IV-A Pretraining
We load ResNet50 weights from DINO [^2] https://github.com/facebookresearch/dino [60] and ViT-S/14 and ViT-B/14 weights (without register) from DINOv2333https://github.com/facebookresearch/dinov2 [16], and conduct continual pretraining for 100 epochs. For ViTs, we randomly mask out 20% patches and send the remaining ones to the trainable encoder. Training is distributed in two nodes each with 4 NVIDIA A100 GPUs and takes 7-30 hours for different backbones. More details can be seen in the appendix.